All strictly monotonically decreasing sequences are convergent.



True or false with full explanation

1
Expert's answer
2022-03-15T10:27:01-0400

Consider "x_n=-n,n\geq 1"

The sequence is obviously strictly monotonically decreasing, and

"\underset{n\rightarrow \infty}{\lim}x_n=-\infty"

Thus it is not convergent.

The statement is false.